UT 05 Sep 2021 - Night #4 of Restart




Night Staff:

OSA: S. Allanson

ISp: R. Reyes

ISA: B. Rothberg, R.T. Gatto (Tucson secret Outpost)

Technical Wizard: J. Hill (Tucson)

Mtn Manager: P. Hartley

Sunset: 18:45

Twlight Ends: 20:04

Twilght Starts: 04:33

Sunrise: 05:52

Starting Config: PEPSI PFU

Night Log:

UT 01:07 - bringing up LUCI-1 test dark exposure, init all

luci1.20210905.0001-0002 - 2.51sec dark, init OK

UT 01:10 - waking up MODS-2, test snaps

-all look OK

-holding off starting up LBCs, unless we get clear skies, unlikely to use them tonight

UT 01:30 - S. Allanson needed to restart GCS for PEPSI again - seems to be a nightly requirement for PEPSI

UT 01:40 - Closed - thick dark clouds overhead - borderline humidity

MODS-2 Focus Checks:

UT 02:12 - starting dualgr2.pro
  • M2B 0003-0005 DON'T USE
  • M2R 0003-0005 DON'T USE
UT 02:25 -took 3 exposure, ln the last M2R IMCS lock failed

-script suggests there should have been more exposures

Redoing dualgr2.pro UT 02:58 - starting redgr2.pro
  • M2R 0017-0027
UT 03:07 - clouds thicker, humidity now at 92% and rising

UT 03:17 - another red IMCS lock failure - retry worked

UT 03:25 - staring dualim2.pro UT 03:43 - another red IMCS lock failure -

-wouldn't let me retry:

STATUS: IMCSLOCK Blue IMCS locked-on, Red Pending BlueIMCS =LockedOn RedIMCS =Pend** ERROR: IMCSLOCK RedIMCS =Fail BlueIMCS =LockedOn
** Abort, Retry, or Ignore? > R

** dualim2.pro aborted with errors at line 70 command 'imcslock'
... aborting IMCS lock request ...
DONE: RED IMCSLOCK Red-Channel IMCS failed to lock-on RedIMCS =Fail

failed at rfoc=540, bfocu=1315

-will re-run from these values to get the last 3 data points

UT 03:54 - starting redim2.pro
  • M2R 0039-0049
UT 04:01 - got some rain or condensation from fog on building (and all-sky)

UT 04:11 - starting blugr2.pro
  • M2B 0028-0038
UT 04:29 - building is soggy now

90% DISK SPACE USED ON MODS2DATA WARNING IT 8061

UT 04:41 - staring bluim2.pro
  • M2B 0039-0049
error in executing script

** ERROR: Command 'instconfig blue imaging' timed out after 90 seconds

running script again

UT 05:06 - running dualpr2.pro UT 05:18 - another red IMCS lock failure, occurred for #0053 - retry worked

UT 05:32 - running redpr2.pro
  • M2R 0061-0067 DO NOT USE
UT 05:44 - another red IMCS lock failure - retry worked

UT 05:45 - another red IMCS lock failure - retry worked

UT 05:46 - another red IMCS lock failure - retry worked

UT 05:47 - another red IMCS lock failure - retry worked

This is now an IMCS lock falure 4 exposures in a row

I think we are in a race condition here - last one reported IMCS lock falure due to an attempt to lock while exposure was reading out

-going to redo this script - i'm not confident of which ones might be affected by the race conditions

UT 05:50 - re-running redpr2.pro

UT 06:53 - instconfig error timed out after 90 seconds - aborted

Calmode timed out

Warning from Log

RGRTILT2 error [2021-09-05T05:54:02] M2.IE> ERROR: RGRTILT2 RGRTILT2=TIMEOUT cannot read from 192.168.139.201:8008

Moved to Gratind Dual first - waited a bit and then re-rean the scripts - so far so good. I think we just got into a bad state - waiting probably helped? IT 6570? if happens again will update

UT 05:59 - redp2.pro
  • M2R 0068-0078
UT 06:13 IMCS timeout before #0075 - retry worked

UT 06:14 - IMCS timeout before #0076 - retry worked -- starting to owrry about the rates...

UT 06:16 - IMCS lock before #0077 - getting into a race condition again...last 3 images usable?

UT 06:20 - running blupr2pro

UT 06:20 - another tiemout for calmode while running the blu2pr2.pro scripts - aborted and retried

BTW I already ran these on UT Sep 02! Updating the master planning wiki page now.

UT 06:31 - Conditions highly unstable so remaining closed for now.

NOTE: I redid Imaging and Grating mode - But PRISM mode is NEW here!

PEPSI PFU AGW checkout...

UT 09:55 - still closed thick clouds, humidity 77%

UT 10:43 - conditions improving - could try to do PEPSI PFU check

UT 10:47 - going for it - opening enclosure - will contact J. Hill to work on the PEPSI checkout

-putting MODS-2 to sleep, no plans to use - plus slightly worried about the RED IMCS issues

-can see a star on SX, DX nothing yet

UT 10:59 - pointing done

UT 11:01 - collination check - R=10 mag star

-will collect WFS images - offline analysis to measure the pupil center

UT 11:05 - collimation complete - seeing~0.74" SX 0.58" DX

UT 11:11 - reconfiguring for bino-LUCI

Field Aberration Collection

  • BS9107 03:32:03.0 +37:20:40.0
  • BS9107_field.list
UT 11:15 - pointing check

UT 11:20 - moving to BS9107 to collimate

UT 11:27 - had to re-check pointing, now collimating again (can actually see the star!)

UT 11:29 - sending GUIDE preset

UT 11:30 - starting field_collect both sides

have to run off OBS2 not ROBS - got a co-pointing error being 25000" away - might have hit wrong list since x-window over ssh slow

BS9107_field.list says "Stone_P" at the top and was attempting to send us off into the ground. So this is a D.Miller error not slow X-windows

Running BS107_new_field.list which contains 16 stars. Stone_P1 had 8 stars (which we did 2 nights ago) - still running on OBS2

Right side is struggling - cant seem to get down to 250nm rms, seeing is <0.7" , lowest we got is 307nm rms.

UT 11:49 - after 21 iterations, moved to first off-axis GS in list (GS23)

-now left is struggling to get to 250nm rms

UT 11:53 -restarting with 350nm rms SX first is #20, DX first is #25

Humidity is 84%

DX is still struggling to reach rms 350nm

Not sure this will be enough stars....

UT 12:20 - done.... we did confirm AGW-2 is now ok.


Proposed Plan of Action:

  • PEPSI
    • PFU Checkout (0.5hr)
      • collimation
      • PFU alignment check for M3
    • PEPSI Pointing Mode - 2hr Medium/High - any conditions
      • Remove retro-reflecting dichroic (to give full FOV)
      • Take images with PFU (Active mode)
      • Need ~28 stars (10th mag or so)
        • PtStars (IRAF) will generate stars based on LST
        • minimized long azimuth slews
      • Analysis with TPOINT - D. Thompson
      • Can do in poor seeing/clouds
  • AGW-2
    • verify cover is flipped and not vignetting - HIGH
    • Field Aberration - field_collect - collect both in parallel
      • Fields:
      • BS9177 18:39:33.7 +49:05:38.0
      • StoneP1 22:37:45.0 -01:13:40.45
      • BS9188 23:30:33.4 +38:18:57.0
      • BS9107 03:32:03.0 +37:20:40.0
  • LBC
    • Standard photometric fields (new ones ready) - requires clear skies (any seeing) - required for ZP calibrations
    • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations
    • Twilight Flats all filters - High (4 x 0.5 hrs - likely min two nights to complete)
    • LBC Pointing Model - (3hr) - any conditions
      • Standard Method
        • <a href='https://wiki.lbto.org/bin/edit/Commissioning/PtStarrs?topicparent=Commissioning.Com2021Restart' rel='nofollow' title='Create this topic'> PtStarrs </a> - please choose targets > 80 deg elevation (more than 1 or 2)
        • Run TMS passively - run 1 reference at the beginning
        • Need stable temperatures
        • Turn on PCS logging
        • Need to collimate at each star
        • Chip 2 only exposure
  • MODS-2
    • MODS Rotator Center measure (use M71 at beg of night or M52 later)
      • requires pointing mask (40" square field stop & dot-circle with radius 45") - high (0.5hr)
    • MODS AGW 5- Measure Rotator zero point (mask alignment) - high (0.5 hr) - mask exchange required
      • M1: dPA =
    • Spectrophotometric standard all modes - medium
      • Grating
      • Prism
    • Test Science OB
      • imaging
      • longslit
  • LUCI-1
    • Field Center Rotation Check (0.5hr) - one field of the three
      • M92 - start
      • M34 - mid
      • M52 - end
    • measure rotator zeropoint (requires MOS mask) - (0.5 hr)
      • ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
    • AGW Transform (if seeing < 1.2")
      • StoneO 21 37 31 +00 31 55 Stone_O_m12.list
      • StoneP 22 35 49 +01 31 33 Stone_P1.list
      • StoneA 00 54 02 -00 32 34 Stone_A3.list
    • Imaging check
      • dither in field (0.5hr) -
      • dither out of field (0.5hr) - M52 K and some J-ban
    • non-sidereal Imaging Check
      • dither in field
      • dither out of field
        • Comet 69/P CG

Current Status

  • LBCs
    • LBC-Blue - available
    • LBC-Red - available
  • MODS
    • MODS-1 - offline - (computers OK, but not ready to pump until Monday)
      • AGW-5 - available
    • MODS-2 - available
      • AGW-6 - available
  • LUCI-1 - available
    • AGW-1 available
  • LUCI-2 offline
    • ´╗┐AGW-2 offline (IT 8457 - Critical)
  • PEPSI - offline Software environmental system issues - earliest Monday
    • AGW-7/8 available


Closed Dome Activities

  • Test Preset before opening (MODS)
  • ARGOS Cal-Unit test for M3 alignment and pupil wobble checkout (alternate to on-sky method)
  • LUCI-1 Only Calibrations
    • Flexure data for G150
      • /home/dthomspons/Luci/FlexureCompensation
      • README - Calib15 - 2hrs
      • Ks filter, Argon lamp, optical sieve mask
      • take dark frame
      • Manual Control MCSGUI for elevations
    • Wavelength data for G150
      • G150 Ks, Argon, take image through slit, then put in grating, line should show up where the slit is located (within a few pixels of the image of the slit)
      • Check a few lines (2 or 3)
  • MODS-2
    • Focus Scripts
    • Bias
      • 2x2
      • 2x1
    • Prism Spectroscopy
      • Dual
      • Red
      • Blue

-- BarryRothberg - 04 Sep 2021
Topic revision: r3 - 05 Sep 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback